a fire hose ejects a stream of wwater at an angle of 35 degree above the horizontal . the water leaves thenozzle with a speed of 25 m /s . assuming that the water behaves like a projectile how far from a building should the fire hose be kept to hit the highest possible fire.

Grade:

1 Answers

xyzabc
12 years ago

at maximum height the horizontal displacement is half the maximum range.

so the range here is 1/2 * u2sin2x/g

or, 1/2 * 625 * 1/20 * sin70 = 1/2 * 625 * 1/20 * 0.77 = 12.03 m answer.